Suppose 10 years ago your mother deposited $3,400 in an investment account earning 5.0% per year. After 3 years she withdrew $1,122. There were no additional deposits or withdrawals from the account. Obtain today's value of the investment account.

$3,710.62

$4,239.39

$4,416.24

$3,959.48
Answer #1

We use the formula:  
A=P(1+r/100)^n
where   
A=future value
P=present value  
r=rate of interest
n=time period

Value after 3 years of deposit

A=3400*(1.05)^3-1122

=2813.925

Hence current value

A=2813.925*(1.05)^7

=$3959.48(Approx)